# Brain network science modelling of sparse neural networks enables Transformers and LLMs to perform as fully connected

### v2

@@ Appendix A Limitations and Future Work
A potential limitation of this work is that the hardware required to accelerate sparse training with unstructured sparsity has not yet become widely adopted. Consequently, this article does not present a direct comparison of training speeds with those of fully connected networks. However, several leading companies [50, 51] have already released devices that support unstructured sparsity in training.
@@ Appendix C Cannistraci-Hebb epitopological rationale
[…]
Here, $u$ and $v$ denote the seed nodes, while $z_{1}$ and $z_{2}$ are common neighbors on the L3 path [32], a walk of three consecutive links that connects $u$ to $v$ via those two intermediate nodes. The term $de^{*}_{i}$ represents the number of external local community links (eLCL) of node $i$, with a default increment of 1 to prevent division by zero. Path-based link prediction has demonstrated its effectiveness on both real-world networks [32] and artificial neural networks [9]. However, this method incurs a high computational cost due to the need to compute and store all length-three paths, resulting in a time complexity of $O(N\cdot d^{3})$, where $N$ is the number of nodes and $d$ is the network’s average degree. This complexity is prohibitive for large models with numerous nodes and higher-density layers. To address this issue, we introduce a more efficient, node-based paradigm that eliminates the reliance on length-three paths between seed nodes. Instead, this approach focuses on the common neighbors of seed nodes. The node-based version of CH3-L3p, denoted as CH2-L3n, is defined as follows:
[…]
Here, $u$ and $v$ denote the seed nodes, while $z$ is the common neighbor on the L3 path [32], a walk of three consecutive links that connects $u$ to $v$ via two of those intermediate nodes. The terms $di^{*}_{z}$ $de^{*}_{z}$ represent the number of internal local community links (iLCLs) and external local community links (eLCLs) of node $i$, with a default increment of 1 to prevent division by zero. Internal local community links (iLCLs) are those that connect nodes belonging to the same local community. Contrarily, external local community links (eLCLs) connect nodes belonging to different communities. Figure 4 gives a visual representation of L2 and L3 paths between two seed nodes $u$ and $v$, defining their local community.
Figure 4: Cannistraci-Hebb epitopological rationale. [32] The figure illustrates an explanatory example of topological link prediction using the Cannistraci-Hebb epitopological rationale based on either L2 or L3 paths. The two black nodes represent the seed nodes whose unobserved interaction is to be assigned a likelihood score. White nodes denote the common neighbours (CNs) of the seed nodes at either L2 or L3 distance. Together, the set of CNs and the internal local community links (iLCL) constitute the local community. Different link types are color-coded: green for nLCLs, red for external local community links (eLCLs), and white for iLCLs. The L2 (path length 2) and L3 (path length 3) paths associated with the illustrated communities are highlighted. Notably, in artificial neural networks (ANNs), linear layers correspond to bipartite networks, which inherently support only L3 path predictions, as shown in Figure 1.
@@ Appendix D Sparse topological initialization
In artificial neural networks (ANNs), fully connected networks are inherently bipartite. This article explores initializing bipartite networks using models from network science. The Bipartite Scale-Free (BSF) [9] network model extends the concept of scale-freeness to bipartite structures, making them suitable for dynamic sparse training. Initially, the BSF model generates a monopartite Barabási-Albert (BA) model [39], a well-established method for creating scale-free networks in which the degree distribution follows a power law ($\gamma$=2.76 in Figure 5). Following the creation of the BA model, the BSF approach removes any connections between nodes of the same type (neuron in the same layer) and rewires these connections to nodes of the opposite type (neuron in the opposite layer). This rewiring is done while maintaining the degree of each node constant to preserve the power-law exponent $\gamma$.
@@
The Bipartite Small-World (BSW) network model [9] is designed to incorporate small-world properties and a high clustering coefficient into bipartite networks. Initially, the model constructs a regular ring lattice and assigns two distinct types of nodes to it. Each node is connected by an equal number of links to the nearest nodes of the opposite type, fostering high clustering but lacking the small-world property. Similar to the Watts-Strogatz model (WS) [38], the BSW model introduces a rewiring parameter, $\beta$, which represents the percentage of links randomly removed and then rewired within the network. At $\beta=1$, the model transitions into an Erdős-Rényi model [52], exhibiting small-world properties but without a high clustering coefficient, which is popular as the topological initialization of the other dynamic sparse training methods [5, 7, 8].
Figure 5: The adjacency matrix of the Bipartite Scale-Free (BSF) network model compared to those of the Bipartite Small-World (BSW) network, the Bipartite Receptive Field with fixed sampling (BRFf), and the Bipartite Receptive field with uniform sampling (BRFu) as parameters $\beta$ and $r$ vary between 0 and 1. a) The BSF model inherently forms a scale-free network characterized by a power-law distribution with $\gamma=2.76$. b) As $\beta$ changes from 0 to 1, the network exhibits reduced clustering. It is important to note that when $\beta=0$, the BSW model does not qualify as a small-world network. c) As $r$ increases towards $1$, the adjacency matrix becomes more random, while sampling the output neurons’ degrees from a fixed or uniform distribution.
@@
The Bipartite Receptive Field (BRF) model is a random network generation technique designed to mimic the receptive field phenomenon in the brain networks. The process involves adding links to the adjacency matrix of the bipartite network, with the connectivity structured around the main diagonal according to a parameter $r\in[0,1]$. A low value of $r$ results in links that are primarily clustered around the diagonal, while a higher value of $r$ leads to a more random connectivity pattern. Specifically, a bipartite adjacency matrix with links near the diagonal indicates that adjacent nodes from the two layers are linked, whereas links far from the diagonal correspond to more distant node pairs.  Mathematically, consider an $N\times M$ bipartite adjacency matrix ${M_{i,j}}_{i=1,\dots,M,j=1,\dots,N}$, where $M$ represents the input size and $N$ represents the output size. Each entry of the matrix $m_{i,j}$ is set to $1$ if input node $i$ is connected to output node $j$, and 00 otherwise. A scoring function $S_{i,j}$ is assigned to each connection in the adjacency matrix based on its distance to the main diagonal. This score is given by:
@@ Appendix E Equal Partition and Neuron Resorting to enhance bipartite scale-free network initialization
As indicated in SET and CHT [5, 9], trained sparse models typically converge to a scale-free network. This suggests that initiating the network with a scale-free structure might initially enhance performance. However, starting directly with a Bipartite Scale-Free model (BSF, power-law exponent $\gamma=2.76$) does not yield effective results. Upon deeper examination, two potential reasons emerge:
@@
  • As demonstrated in CHT, in the final network, the hub nodes of one layer’s output should correspond to the input layer of the subsequent layer, which means the hub nodes should have a high degree on both sides of the layer. However, the BSF model’s random selection disrupts this correspondence, significantly reducing the number of Credit Assignment Paths (CAP) [9] in the model. CAP is defined as the chain of transformation from input to output, which counts the number of links that go through the hub nodes in the middle layers.
@@
We have adapted network percolation [53, 9] to suit the architecture of the Transformer after link removal. The core idea is to identify inactive neurons, which are characterized by having no connections on either one or both sides within a layer of neurons. Such neurons disrupt the flow of information during forward propagation or backpropagation. In addition, Layer-wise computation of the CH link prediction score further implies that neurons without connections on one side are unlikely to form connections in the future. Therefore, network percolation becomes essential to optimize the use of remaining links.
As shown in Figure 1, network percolation encompasses two primary processes: c1) inactive neuron removal to remove the neurons that lack connections on one or both sides; c2) incomplete path adjustment to remove the incomplete paths where links connect to the inactive neurons after c1). Typically applied in simpler continuous layers like those in an MLP, network percolation requires modification for more complex structures. For example, within the Transformer’s self-attention module, the outputs of the query and key layers undergo a dot product operation. It necessitates percolation in these layers to examine the activity of the neurons in both output layers at the same position. Similar interventions are necessary in the up_proj and gate_proj layers of the MLP module in the LLaMA model family [1, 54].
@@ G.1 Fixed Density Dynamic Sparse Training Methods
[5]: Removes connections based on weight magnitude and randomly regrows new links.
@@
[7]: Removes connections based on weight magnitude and regrows links using gradient information, gradually reducing the proportion of updated connections over time.
@@
[9]: A state-of-the-art (SOTA) gradient-free method that removes links with weight magnitude and regrows links based on CH3-L3 scores. CHT is often applied with early stopping to mitigate its computational complexity when working with large models.
@@ G.2 Gradual Density Decrease Dynamic Sparse Training Methods
[55, 43]: Prunes the network with weight magnitude and gradually decreases the density based on Equation 10. Although originally a pruning method, GMP is treated as a dynamic sparse training method in their implementation [43], as it stores historical weights and allows pruned weights to reappear during training, since, during training, the pruning threshold might change.
@@
[8]: Implements a two-stage density decrease strategy as described in the original work. It removes links based on the combination of weight magnitude and 0.01*gradient and regrows new links randomly.
@@
[42]: Gradually decreases density using Equation 10. Similar to RigL, GraNet removes links based on the weight magnitude and regrows new links with the gradient of the existing links.
@@ Appendix H Ablation and Sensitivity Tests
For sparse topological initialization, we compare BRF, BSW, BSF, and CSTI [9] across three image classification datasets, as shown in Figure 8b. The results indicate that when the inputs can directly access task-relevant information, CSTI consistently achieves the best performance. In general, BRF and BSW perform similarly under these conditions, but outperform the BSF initialization.
@@
We first conduct a simple evaluation of the link removal methods introduced in this article when changing the $\alpha$ and $\delta$ inside Figure 1b2) on Figure 6b) and Figure 7b). The removal methods are selected from Weight Magnitude (WM), Weight Magnitude soft (WMs), Relative Importance (RI), and Relative Importance soft (RIs). For WM we fix the hyperparameters $\alpha=1$ and $\delta=1$; for RI we fix the hyperparameters $\alpha=0$ and $\delta=0.5$; for WMs we fix $\alpha=1$ and we let $\delta$ increase linearly from $0.5$ to $0.9$; for RIs we fix $\alpha=0$ and let $\delta$ increase linearly from $0.5$ to $0.9$. From the results, it can be observed that WMs performs the best in most cases. We compare these methods with those in [56] in Table 13 on two machine translation tasks. The results indicate that using WMs as a link removal method generally outperforms the alternatives.

### v3

@@ Appendix A Limitations and Future Work
A potential limitation of this work is that the hardware required to accelerate sparse training with unstructured sparsity has not yet become widely adopted. Consequently, this article does not present a direct comparison of training speeds with those of fully connected networks. However, several leading companies [thangarasa2023spdf, pmlr-v119-kurtz20a] have already released devices that support unstructured sparsity in training.
@@ Appendix C Cannistraci-Hebb epitopological rationale
$\textbf{CH3-L3p}(u,v)=\sum_{z_{1},z_{2}\in L3}\frac{1}{\sqrt{de^{*}_{z_{1}}\cdot de^{*}_{z_{2}}}}$ (5)
Here, $u$ and $v$ denote the seed nodes, while $z_{1}$ and $z_{2}$ are common neighbors on the L3 path [CHA], a walk of three consecutive links that connects $u$ to $v$ via those two intermediate nodes. The term $de^{*}_{i}$ represents the number of external local community links (eLCL) of node $i$, with a default increment of 1 to prevent division by zero. Path-based link prediction has demonstrated its effectiveness on both real-world networks [CHA] and artificial neural networks [zhang2023ultra]. However, this method incurs a high computational cost due to the need to compute and store all length-three paths, resulting in a time complexity of $O(N\cdot d^{3})$, where $N$ is the number of nodes and $d$ is the network’s average degree. This complexity is prohibitive for large models with numerous nodes and higher-density layers. To address this issue, we introduce a more efficient, node-based paradigm that eliminates the reliance on length-three paths between seed nodes. Instead, this approach focuses on the common neighbors of seed nodes. The node-based version of CH3-L3p, denoted as CH2-L3n, is defined as follows:
$\textbf{CH2-L3n}(u,v)=\sum_{z\in L3}\frac{di_{z}^{*}}{de^{*}_{z}}$ (6)
Here, $u$ and $v$ denote the seed nodes, while $z$ is the common neighbor on the L3 path [CHA], a walk of three consecutive links that connects $u$ to $v$ via two of those intermediate nodes. The terms $di^{*}_{z}$ $de^{*}_{z}$ represent the number of internal local community links (iLCLs) and external local community links (eLCLs) of node $i$, with a default increment of 1 to prevent division by zero. Internal local community links (iLCLs) are those that connect nodes belonging to the same local community. Contrarily, external local community links (eLCLs) connect nodes belonging to different communities. Figure 4 gives a visual representation of L2 and L3 paths between two seed nodes $u$ and $v$, defining their local community.
Figure 4: Cannistraci-Hebb epitopological rationale. [CHA] The figure illustrates an explanatory example of topological link prediction using the Cannistraci-Hebb epitopological rationale based on either L2 or L3 paths. The two black nodes represent the seed nodes whose unobserved interaction is to be assigned a likelihood score. White nodes denote the common neighbours (CNs) of the seed nodes at either L2 or L3 distance. Together, the set of CNs and the internal local community links (iLCL) constitute the local community. Different link types are color-coded: green for nLCLs, red for external local community links (eLCLs), and white for iLCLs. The L2 (path length 2) and L3 (path length 3) paths associated with the illustrated communities are highlighted. Notably, in artificial neural networks (ANNs), linear layers correspond to bipartite networks, which inherently support only L3 path predictions, as shown in Figure 1.
@@ Appendix D Sparse topological initialization
In artificial neural networks (ANNs), fully connected networks are inherently bipartite. This article explores initializing bipartite networks using models from network science. The Bipartite Scale-Free (BSF) [zhang2023ultra] network model extends the concept of scale-freeness to bipartite structures, making them suitable for dynamic sparse training. Initially, the BSF model generates a monopartite Barabási-Albert (BA) model [barabasi1999emergence], a well-established method for creating scale-free networks in which the degree distribution follows a power law ($\gamma$=2.76 in Figure 5). Following the creation of the BA model, the BSF approach removes any connections between nodes of the same type (neuron in the same layer) and rewires these connections to nodes of the opposite type (neuron in the opposite layer). This rewiring is done while maintaining the degree of each node constant to preserve the power-law exponent $\gamma$.
@@
The Bipartite Small-World (BSW) network model [zhang2023ultra] is designed to incorporate small-world properties and a high clustering coefficient into bipartite networks. Initially, the model constructs a regular ring lattice and assigns two distinct types of nodes to it. Each node is connected by an equal number of links to the nearest nodes of the opposite type, fostering high clustering but lacking the small-world property. Similar to the Watts-Strogatz model (WS) [watts1998collective], the BSW model introduces a rewiring parameter, $\beta$, which represents the percentage of links randomly removed and then rewired within the network. At $\beta=1$, the model transitions into an Erdős-Rényi model [erdds1959random], exhibiting small-world properties but without a high clustering coefficient, which is popular as the topological initialization of the other dynamic sparse training methods [SET, RIGL, MEST].
Figure 5: The adjacency matrix of the Bipartite Scale-Free (BSF) network model compared to those of the Bipartite Small-World (BSW) network, the Bipartite Receptive Field with fixed sampling (BRFf), and the Bipartite Receptive field with uniform sampling (BRFu) as parameters $\beta$ and $r$ vary between 0 and 1. a) The BSF model inherently forms a scale-free network characterized by a power-law distribution with $\gamma=2.76$. b) As $\beta$ changes from 0 to 1, the network exhibits reduced clustering. It is important to note that when $\beta=0$, the BSW model does not qualify as a small-world network. c) As $r$ increases towards 11, the adjacency matrix becomes more random, while sampling the output neurons’ degrees from a fixed or uniform distribution.
@@
The Bipartite Receptive Field (BRF) model is a random network generation technique designed to mimic the receptive field phenomenon in the brain networks. The process involves adding links to the adjacency matrix of the bipartite network, with the connectivity structured around the main diagonal according to a parameter $r\in[0,1]$. A low value of $r$ results in links that are primarily clustered around the diagonal, while a higher value of $r$ leads to a more random connectivity pattern. Specifically, a bipartite adjacency matrix with links near the diagonal indicates that adjacent nodes from the two layers are linked, whereas links far from the diagonal correspond to more distant node pairs. Mathematically, consider an $N\times M$ bipartite adjacency matrix ${M_{i,j}}_{i=1,\dots,M,j=1,\dots,N}$, where $M$ represents the input size and $N$ represents the output size. Each entry of the matrix $m_{i,j}$ is set to $1$ if input node $i$ is connected to output node $j$, and 0 otherwise. A scoring function $S_{i,j}$ is assigned to each connection in the adjacency matrix based on its distance to the main diagonal. This score is given by:
@@ Appendix E Equal Partition and Neuron Resorting to enhance bipartite scale-free network initialization
As indicated in SET and CHT [SET, zhang2023ultra], trained sparse models typically converge to a scale-free network. This suggests that initiating the network with a scale-free structure might initially enhance performance. However, starting directly with a Bipartite Scale-Free model (BSF, power-law exponent $\gamma=2.76$) does not yield effective results. Upon deeper examination, two potential reasons emerge:
@@
  • As demonstrated in CHT, in the final network, the hub nodes of one layer’s output should correspond to the input layer of the subsequent layer, which means the hub nodes should have a high degree on both sides of the layer. However, the BSF model’s random selection disrupts this correspondence, significantly reducing the number of Credit Assignment Paths (CAP) [zhang2023ultra] in the model. CAP is defined as the chain of transformation from input to output, which counts the number of links that go through the hub nodes in the middle layers.
@@
We have adapted network percolation [percolation, zhang2023ultra] to suit the architecture of the Transformer after link removal. The core idea is to identify inactive neurons, which are characterized by having no connections on either one or both sides within a layer of neurons. Such neurons disrupt the flow of information during forward propagation or backpropagation. In addition, Layer-wise computation of the CH link prediction score further implies that neurons without connections on one side are unlikely to form connections in the future. Therefore, network percolation becomes essential to optimize the use of remaining links.
As shown in Figure 1, network percolation encompasses two primary processes: c1) inactive neuron removal to remove the neurons that lack connections on one or both sides; c2) incomplete path adjustment to remove the incomplete paths where links connect to the inactive neurons after c1). Typically applied in simpler continuous layers like those in an MLP, network percolation requires modification for more complex structures. For example, within the Transformer’s self-attention module, the outputs of the query and key layers undergo a dot product operation. It necessitates percolation in these layers to examine the activity of the neurons in both output layers at the same position. Similar interventions are necessary in the up_proj and gate_proj layers of the MLP module in the LLaMA model family [touvron2023llama, touvron2023llama2].
@@ G.1 Fixed Density Dynamic Sparse Training Methods
[SET]: Removes connections based on weight magnitude and randomly regrows new links.
@@
[RIGL]: Removes connections based on weight magnitude and regrows links using gradient information, gradually reducing the proportion of updated connections over time.
@@
[zhang2023ultra]: A state-of-the-art (SOTA) gradient-free method that removes links with weight magnitude and regrows links based on CH3-L3 scores. CHT is often applied with early stopping to mitigate its computational complexity when working with large models.
@@ G.2 Gradual Density Decrease Dynamic Sparse Training Methods
[han2015gmp, zhu2017gmp]: Prunes the network with weight magnitude and gradually decreases the density based on Equation 10. Although originally a pruning method, GMP is treated as a dynamic sparse training method in their implementation [zhu2017gmp], as it stores historical weights and allows pruned weights to reappear during training, since, during training, the pruning threshold might change.
@@
[MEST]: Implements a two-stage density decrease strategy as described in the original work. It removes links based on the combination of weight magnitude and 0.01*gradient and regrows new links randomly.
@@
[liu2021granet]: Gradually decreases density using Equation 10. Similar to RigL, GraNet removes links based on the weight magnitude and regrows new links with the gradient of the existing links.
@@ Appendix H Ablation and Sensitivity Tests
For sparse topological initialization, we compare BRF, BSW, BSF, and CSTI [zhang2023ultra] across three image classification datasets, as shown in Figure 8b. The results indicate that when the inputs can directly access task-relevant information, CSTI consistently achieves the best performance. In general, BRF and BSW perform similarly under these conditions, but outperform the BSF initialization.
@@
We first conduct a simple evaluation of the link removal methods introduced in this article when changing the $\alpha$ and $\delta$ inside Figure 1b2) on Figure 6b) and Figure 7b). The removal methods are selected from Weight Magnitude (WM), Weight Magnitude soft (WMs), Relative Importance (RI), and Relative Importance soft (RIs). For WM we fix the hyperparameters $\alpha=1$ and $\delta=1$; for RI we fix the hyperparameters $\alpha=0$ and $\delta=0.5$; for WMs we fix $\alpha=1$ and we let $\delta$ increase linearly from $0.5$ to $0.9$; for RIs we fix $\alpha=0$ and let $\delta$ increase linearly from $0.5$ to $0.9$. From the results, it can be observed that WMs performs the best in most cases. We compare these methods with those in [nowak2023fantasticweightsthemprune] in Table 12 on two machine translation tasks. The results indicate that using WMs as a link removal method generally outperforms the alternatives.
@@
Density Decay Strategy.
To further demonstrate the benefit of the density decay strategy, we conduct an ablation study comparing the sigmoid-based density decay strategy with a cubic-based density decay strategy. We evaluate the performance on the LLaMA-60M model using the C4 dataset and report the perplexity. We show the number of density decay steps and a comparison between the cubic density decay that GraNet [liu2021granet] and GMP [zhu2017gmp] introduced and our proposed sigmoid density decay strategy. The results in the Table 13 show that the sigmoid-based decay consistently achieves lower PPL than the cubic-based decay in all the density decay steps, and both strategies outperform the baseline fixed sparsity (CHTs). These results confirm the motivation for introducing the sigmoid density decay strategy. The table below shows the PPL of CHTss with the two density decay strategies over different density decay steps.
Table 13: Comparison of CHTss variants (Sigmoid- vs. Cubic-based) under different density decay steps. Baseline (CHTs) achieves 35.59. Lower values indicate better performance.
Density Decay Steps 1000 2000 3000 4000 5000 6000 7000 8000 9000
CHTss (Sigmoid-based) 35.20 35.19 35.24 34.63 34.68 34.63 34.46 34.50 34.39
CHTss (Cubic-based) 35.23 35.21 35.29 35.03 34.96 39.23 34.96 34.79 34.94
Curvature of the Density Decay.
In CHTss, the parameter $k$ dictates the pruning schedule’s shape: lower $k$ values yield a smoother decay, while higher values create a sharper curve. We varied $k$ in 2, 4, 6, 8, 10 on LLaMA60M at 70% and 95% sparsity, using perplexity as the evaluation metric. Results on Table 14 show that overly sharp decay curves (larger $k$) cause instability. In the article, we used $k=6$ for all experiments.
Table 14: Perplexity comparison of LLaMA60M under different initial sparsity and $k$ values at 70% and 95% sparsity levels. Lower is better.
LLaMA60M (70% sparsity) k=2.0k{=}2.0 k=4.0k{=}4.0 k=6.0k{=}6.0 k=8.0k{=}8.0 k=10.0k{=}10.0
0.1 27.61 27.74 27.74 27.82 29.79
0.2 27.60 28.81 27.81 29.82 34.05
0.3 27.66 27.67 27.78 27.59 27.58
0.4 27.93 27.82 27.78 27.73 27.71
0.5 28.05 27.82 27.62 27.83 28.13
LLaMA60M (95% sparsity) k=2.0k{=}2.0 k=4.0k{=}4.0 k=6.0k{=}6.0 k=8.0k{=}8.0 k=10.0k{=}10.0
0.1 37.00 36.02 35.86 36.42 36.59
0.2 36.83 36.08 35.42 38.39 35.87
0.3 36.67 35.30 35.53 35.67 35.78
0.4 35.71 35.48 35.42 35.44 35.63
0.5 36.76 35.70 35.59 35.72 36.05
Removal Fraction $\zeta$.
This parameter determines the proportion of existing connections that are pruned during each pruning-regrowth cycle. It directly controls the amount of structural change introduced to the network at each step. In our sensitivity analysis on Table 15, we evaluated a range of $\zeta$ values (0.1, 0.2, 0.3, 0.4, 0.5) to assess their effect on the convergence and final performance of the CHTs model. The tests were performed on MLPs (CIFAR-10) at 90%, 95%, and 99% sparsities. Results are averaged over three seeds. When the sparsity is lower, we need a lower $\zeta$ to keep the model trained more stably, and when the sparsity is higher, a higher $\zeta$ is required to encourage the topological exploration. In all experiments involving MLP and Transformer, we use a $\zeta$ of 0.3, and for LLaMA models, we use 0.1 for all the experiments.
Table 15: Test accuracy (%) on CIFAR-10 under varying sparsity ratios and $\zeta$ values. Mean $\pm$ standard deviation over multiple runs. Bold indicates best per column.
$\zeta$ CIFAR-10 (90%) CIFAR-10 (95%) CIFAR-10 (99%)
0.1 68.73 $\pm$ 0.27 69.28 $\pm$ 0.08 63.45 $\pm$ 0.04
0.2 68.63 $\pm$ 0.10 69.78 $\pm$ 0.09 67.30 $\pm$ 0.21
0.3 68.47 $\pm$ 0.10 69.37 $\pm$ 0.10 68.08 $\pm$ 0.02
0.4 67.90 $\pm$ 0.10 69.01 $\pm$ 0.08 67.83 $\pm$ 0.19
0.5 67.47 $\pm$ 0.14 68.56 $\pm$ 0.15 67.24 $\pm$ 0.14
CHTs vs. CHTss.
In most experiments, CHTss (with the sigmod density decay schedule) consistently outperforms CHTs. There are, however, a few counter-examples (e.g., LLaMA-1B at 70% target sparsity on OpenWebText) where CHTs slightly surpasses CHTss. To understand this outlier behavior, we first note that in the LLaMA-1B OpenWebText run the schedule only decreased density from 50% to 30%—a small change relative to the fixed 50% baseline—yielding very similar perplexities (14.66 vs. 15.15). When we extended the schedule to a much lower final density 5%, the advantage of CHTss became clear: CHTss reached 16.51 PPL versus 18.93 for CHTs. This aligns with our itop analysis: at 30% density, over 90% of candidate links appear during the evolution of both methods, but at 5% density CHTs touches only $\sim$20% of links while CHTss still explores $\sim$90%. In other words, density–decay confers a broader search over plausible connections precisely where search matters most—at high sparsity. We further corroborated this with a cross-scale study on six LLaMA sizes (20M($\rightarrow$)1B) at 70%, 90% , and 95% sparsity on Table 16: averaged validation perplexity differences are not significant at 70% (Wilcoxon (p=0.688)), but become significant in favor of CHTss at 90% and 95% (p=0.031 for both). Taken together, these results indicate that occasional wins of CHTs arise in regimes where the decay is shallow and the effective search space is already well covered; they do not suggest that the sigmoid density–decay strategy fails on larger models or particular tasks. Rather, its benefit scales with the target sparsity: the more aggressive the final sparsity, the more CHTss helps by exposing and testing a larger set of candidate links during training.
Table 16: CHTs vs. CHTss across sparsity levels (validation perplexity; lower is better).
Sparsity Mean CHTs $\downarrow$ Mean CHTss $\downarrow$ $p$-value Significance
0.70 23.81 23.88 0.688 Not significant
0.90 28.47 26.51 0.031 Significant
0.95 31.89 29.19 0.031 Significant